Bob's age pluses 3is a square number. Bob's age minus 3 isBob's age plus 3. What is Bob's age?
Answer by Alan3354(69443)

Bob's age minus 3 isBob's age plus 3. What is Bob's age?
-------------
B - 3 = B + 3
-3 = 3
No solution.